Solve (y ^ 2 * z)/x * p + xzq = y ^ 2 .

1
Expert's answer
2022-02-08T14:30:37-0500

"y^2zp" "+x^2zq=y^2x"


The subsidiary equations are

"\frac{dx}{y^2z}=\frac{dy}{x^2z}=\frac{dz}{y^2x}"


Considering the first two terms of the equation we get

"\frac{dx}{y^2}=\frac{dy}{x^2}"


"x^2dx-y^2dy=0"

"x^3-y^3=c_1"


Considering the 1st and last terms of the equation

"\frac{dx}{y^2z}=\frac{dz}{y^2x}"


"\frac{dx}{z}=\frac{dz}{x}"


"xdx-zdz=0"

"x^2-z^2=c_2"


Solution


"f(x^3-y^3,x^2-z^2)=0"